For each of the following situations give the degrees of freedom and an appropriate bound on the P-value (give the exact value if you have software available) for the χ2 statistic for testing the null hypothesis of no association between the row and column variables.

(a) A 2 by 2 table with χ2 = 0.94.

df = 1
P-value = 2


(b) A 4 by 4 table with χ2 = 17.32.

df = 3
P-value = 4


(c) A 2 by 8 table with χ2 = 22.47.

df = 5
P-value = 6


(d) A 5 by 3 table with χ2 = 13.26.

df = 7
P-value = 8
Answer #1

(a). For a 2 X 2 table with chi ^{2} = 0.94,

(1). the value of  the degrees of freedom is = (k-1)(l-1) = (2-1)×(2-1)=1×1=1 .

(2). The P-value of chi ^{2} test is Pl)->= X2) = 0.332278 , at 0.05 level of significance.

(b). For a 4×4 table with chi ^{2} = 17.32,

(3). The value of  the degrees of freedom is = (k-1)(l-1) =(4-1)×(4-1)=3×3=9

(4). The P-value of chi ^{2} test is Pl)->= X2) = 0.043934 , at 0.05 level of significance

(c). For a 2 X 8 table with chi ^{2} = 22.47,

(5). ​The value of  the degrees of freedom is = (k-1)(l-1) =(2-1)×(8-1)=1×7=7

(6). The P-value of chi ^{2} test is P(X >= X2) 0.002107 , at 0.05 level of significance

(d). For a 5 imes 3 table with chi ^{2} = 13.26,

  (7). The value of  the degrees of freedom is = (k-1)(l-1) =(5-1)×(3-1)=4×2=8

(8). The P-value of chi ^{2} test is PlX >= X2) = 0.103211 , at 0.05 level of significance